### v1

@@ 4 Phase noise
In this section, we report the phase noise of the signal from the TDO operating at 11 mK through port AC I using a spectrum analyzer (R&S FSV3030). Additionally, we sent the signal from the commercial microwave source through port AC II and measured the phase noise of the reflected signal through port AC I. The results are compared and shown in Fig. 4. Note that the phase noise of the commercial microwave source, measured directly on the spectrum analyzer without routing it to the device at low temperature, showed identical results. This indicates that the measured phase noise originates from the microwave sources and not from the device at low temperature.
We found that using a lead-acid battery to bias the TDO instead of the Yokogawa GS200 improves the phase noise as shown in Fig. 4 (see also Sec. C for details). The measured phase noise of the TDO, powered by a battery, was $-115$ dBc/Hz at a 1 MHz offset, demonstrating comparable or superior performance to CMOS devices operating at 4 K in a similar carrier frequency range [33, 34, 11]. While CMOS devices achieve better phase noise in the lower- and higher-frequency ranges, the TDO provides better performance within the mid-frequency range. Moreover, with the battery as its power source, the TDO’s phase noise is either comparable to or outperforms that of a commercial microwave source (Vaunix Lab-Brick LMS-451D) across the entire measured frequency range.
Figure 4: Measured phase noise relative to the carrier frequency (around 141.8 MHz) as a function of offset frequency with a bandwidth of 1 Hz. The phase noise of four different commercial microwave sources is shown: Rigol DSA815 (black), Analog Devices ADF4351 (blue), Vaunix Lab-Brick LMS-451D (red), along with the TDO powered by Yokogawa GS200 DC source (green) and by a lead-acid battery (light green). For comparison, the phase noise generated by a CMOS device at a carrier frequency of 600 MHz is also shown (orange) [33].
@@ 5 Amplitude stability
In this section, we measured the amplitude of the signal from the TDO operating at 11 mK through port AC I. The TDO generated a signal at 141.8 MHz with $V_{\mathrm{TD}}=0.18$ V and $V_{\mathrm{VD}}=-0.5$ V. The signal, recorded at a sampling rate of 20 GS/s for 1.6 ms, was passed through a digital bandpass filter centered at the carrier frequency with a 50 MHz bandwidth (see Appendix D). The amplitude was then extracted using a Hilbert transform and is plotted in Fig. 5(a) and its histogram is shown in Fig. 5(b). Similar to the phase noise measurement, the signal from a commercial microwave source was sent through port AC II, and the amplitude stability of the reflected signal measured through port AC I is also shown for comparison in Fig. 5. For validation of our analytical method, a synthetic signal at 141.8 MHz generated using a sine function is also plotted. Regardless of whether a lead-acid battery or the Yokogawa GS200 was used as the DC source, the amplitude stability of the TDO was superior to that of any commercial microwave source used in this study.
@@ 6 Discussion
We do not yet know the origin of the amplitude fluctuations, and identifying it will be the subject of future work. Since the oscilloscope used in this study has an 8-bit resolution, its resolution of $1/2^{8}=0.39\%$ is comparable to the observed fluctuations of $45\,\mu\text{V}/15\,\text{mV}=0.3\%$. This suggests that the fluctuations might arise from the resolution limit of the oscilloscope rather than being intrinsic to the TDO. To investigate this further, improvements in the measurement setup, such as using higher-resolution ADCs, will be necessary. For qubit readout, particularly for electrons in semiconductors being used as qubits, where the qubit state is determined by observing the amplitude distribution [14, 15, 16], it is unlikely that the TDO we developed would result in worse readout fidelity compared to the commercial microwave sources used here. However, looking forward, it will be necessary to evaluate the level of amplitude stability required for qubit readout using actual qubits.
@@
Although the current study focuses on qubit readout, the TDO could potentially be utilized for qubit operation although its phase noise is currently not superior to that of commercial sources and could become a limiting factor if used for qubit operation [35]. Further studies are needed to determine the phase noise requirements for qubit operation. Moreover, for qubit operation, the frequency of 140 MHz is too low. For operations involving electron spin states, GHz-range frequencies would be required. To achieve oscillation at higher frequencies, it is essential to reduce the parasitic capacitance of the tunnel diode and the circuit board. The parasitic capacitance of tunnel diodes exhibits significant variation, which is unlikely to be solely due to the pn-junction capacitance. Instead, this variation may originate from the diode packaging. Therefore, rather than relying on commercially available tunnel diodes, designing a tunnel diode specifically optimized for qubit readout or operation applications would be more effective. Minimizing the parasitic capacitance on the circuit board is also important. One possible approach is to fabricate the entire circuit board using microfabrication techniques. Another approach is to replace bulky varactor diodes with variable capacitors made from ferroelectric materials, such as strontium titanate (STO) [36]. Ferroelectric materials like STO not only allow for variable capacitance but also eliminate leakage currents, reducing circuit losses and providing a significant advantage over traditional varactor diodes.
@@
Considering that the cooling power of a typical cryogenic refrigerator’s 4 K stage is approximately 200 mW, and its mixing chamber stage provides about $400\,\mu\text{W}$ at 100 mK, it is feasible to place up to 20,000 microwave sources on the 4 K stage and 400 on the mixing chamber stage. These numbers are promising for realizing a scalable quantum computer; nevertheless, further improvements are necessary. In principle, lower power consumption of tunnel diodes can be fabricated by making the surface area of the tunnel diode smaller. While this would also lead to a reduction in output power, it can be compensated by adjusting the ratio of the number of turns in the spiral inductor to the number of turns in the pickup coil. However, the impact of back action on the oscillator must be carefully considered to ensure reliable operation.
Considering the need to apply a magnetic field for electron spin qubits, improvements such as using magnetic field-resilient superconducting materials like NbTiN instead of Nb and designing spiral inductors with these materials could be explored.
@@ 7 Conclusion
In conclusion, we developed a 140 MHz tunnel diode oscillator with low power consumption of 1 $\mu$W and characterized its performance at 11 mK. The achieved output power and frequency fall within the range typically used for qubit readout for electrons in semiconductors being used as qubits. The amplitude stability is better than that of the commercial microwave sources measured in this study, which is encouraging for its application in qubit readout.
@@ Acknowledgements
Appendix A Temperature dependence
@@
The sensitivity was found to be 0.3 kHz/mK in the temperature range of 60 mK to 120 mK, with no detectable dependence below 60 mK with a slight hysteresis. This result is significantly better than the temperature dependence reported in previous studies above 1 K  [8].
The observed temperature dependence between 60 mK and 120 mK is unlikely to be due to changes in the tunnel diode’s I-V curve, as no significant variation in the I-V characteristics is observed in this temperature range, making it improbable that frequency shifts result from a change in the operating point.
In prior studies measuring above 1 K [8], temperature dependence was attributed to the surface impedance, thermal expansion, and magnetic susceptibility induced by iron impurities in the copper used in the inductor. In contrast, this work uses a superconducting material, Nb, with a critical temperature $T_{c}$ of 6 K, in the inductor. Given that the measured temperature range is significantly below $T_{c}$, it is unlikely that the temperature dependence is primarily due to changes in the properties of Nb, though some contributions cannot be ruled out. Alternatively, the temperature dependence might arise from changes in the physical properties of the metals used in the PCB substrate.
Appendix B Varactor diode IV curve
@@
Appendix C DC voltage source stability
@@
Appendix D Time-domain data
@@
Figure 8: Time-domain data measured via port AC II using a 20 GS/s sampling rate oscilloscope (LeCroy WaveRunner 9054). The data include signals from the TDO powered by Yokogawa GS200 DC source (green) and a lead-acid battery (light green), as well as four different commercial microwave sources: Rigol DSA815 (black), Analog Devices ADF4351 (blue), and Vaunix Lab-Brick LMS-451D (red). The respective carrier frequencies are $f_{0}=$141.800 MHz, 141.841 MHz, 141.802 MHz, 141.806 MHz, and 141.803 MHz. For the TDO, $V_{\mathrm{TD}}=0.18$ V, and $V_{\mathrm{VD}}=-0.5$ V. For a consistency check, a synthetic signal at 141.8 MHz with 15 mV amplitude, generated using a sine function, is also plotted. The raw signal is shown in (a), while (b) presents the signal after applying a digital bandpass filter with a 50 MHz bandwidth centered on the carrier frequency.
Appendix E Power spectra
@@
where the time-domain data $x_{n}$ is shown in Fig. D(b) and $X_{k}$ represents the Fourier coefficients. The root-mean-square amplitude in volts was calculated as:

### v2

@@ 4 Phase noise
In this section, we report the phase noise of the signal from the TDO operating at MCT=11 mK through AC port I using a spectrum analyzer (R&S FSV3030). Additionally, we sent the signal from the commercial microwave source through AC port II and measured the phase noise of the reflected signal through AC port I. The results are compared and shown in Fig. 4. Note that the phase noise of the commercial microwave source, measured directly on the spectrum analyzer without routing it to the device at low temperature, showed almost identical results. This indicates that the measured phase noise originates from the microwave sources and not from the device at low temperature.
We found that using a lead-acid battery to bias the TDO instead of the Yokogawa GS200 improves the phase noise as shown in Fig. 4 (see also Sec. D for details). The measured phase noise of the TDO, powered by a battery, was $-115$ dBc/Hz at a 1 MHz offset, demonstrating comparable or superior performance to CMOS devices operating at the 4 K stage in a similar carrier frequency range [35, 36, 11]. While CMOS devices achieve better phase noise in the lower- and higher-frequency ranges, the TDO provides better performance within the mid-frequency range. Moreover, with the battery as its power source, the TDO’s phase noise is either comparable to or outperforms that of a commercial microwave source (Vaunix Lab-Brick LMS-451D) across the entire measured frequency range.
The TDO has comparable or higher phase noise in the low-frequency range compared to commercial microwave sources. To improve this, enhancing the stability of the DC bias could help mitigate noise fluctuations. Additionally, the phase noise might be caused by two-level systems in the device, suggesting that improvements in the tunnel diode itself may be necessary. Another potential improvement is integrating a feedback mechanism to actively stabilize the frequency or implementing advanced filtering techniques to suppress unwanted noise components.
Figure 4: Measured phase noise relative to the carrier frequency (around 141.8 MHz) as a function of offset frequency with a bandwidth of 1 Hz. The phase noise of four different commercial microwave sources is shown: Rigol DSA815 (black), Analog Devices ADF4351 (blue), Vaunix Lab-Brick LMS-451D (red), along with the TDO powered by Yokogawa GS200 DC source (green) and by a lead-acid battery (light green). For comparison, the phase noise generated by a CMOS device at a carrier frequency of 600 MHz is also shown (orange) [35]. The jumps observed below $2\times 10^{5}$ Hz are likely caused by noise from the pulse tube of the dilution refrigerator. The jumps around $10^{6}$ Hz are attributed to interference from a nearby radio station transmitting at 810 kHz (Appendix D). The jumps above $10^{8}$ Hz are due to harmonics of the carrier frequency.
@@ 5 Amplitude stability
In this section, we measured the amplitude of the signal from the TDO operating at MCT=11 mK through AC port I. The TDO generated a signal at 141.8 MHz with $V_{\mathrm{TD}}=0.18$ V and $V_{\mathrm{VD}}=-0.5$ V. The signal, recorded at a sampling rate of 20 GS/s for 1.6 ms, was passed through a digital bandpass filter centered at the carrier frequency with a 50 MHz bandwidth (see Appendix E). The amplitude was then extracted using a Hilbert transform and is plotted in Fig. 5(a), and its histogram is shown in Fig. 5(b). Similar to the phase noise measurement, a signal from a commercial microwave source was sent through AC port II, and the amplitude stability of the reflected signal, measured via AC port I, is shown in Fig. 5 for comparison. For validation of our analytical method, a synthetic sine signal at 141.8 MHz generated is also plotted. Regardless of whether a lead-acid battery or the Yokogawa GS200 was used as the DC source, the amplitude stability of the TDO was superior to that of any commercial microwave source used in this study.
@@ 6 Discussion
We do not yet know the origin of the amplitude fluctuations, and identifying it will be the subject of future work. Since the oscilloscope used in this study has an 8-bit resolution, its resolution of $1/2^{8}=0.39\%$ is comparable to the observed fluctuations of $45\,\mu\text{V}/15\,\text{mV}=0.3\%$. This suggests that the fluctuations might arise from the resolution limit of the oscilloscope rather than being intrinsic to the TDO. To investigate this further, improvements in the measurement setup, such as using higher-resolution ADCs, will be necessary. For qubit readout, particularly for electrons in semiconductors, where the qubit state is determined by observing the amplitude distribution [14, 15, 16], it is unlikely that the TDO we developed would result in worse readout fidelity compared to the commercial microwave sources used here. However, looking forward, it will be necessary to evaluate the level of amplitude stability required for qubit readout using actual qubits.
@@
Although the current study focuses on qubit readout, the TDO could potentially be utilized for qubit operation. However, its phase noise is currently not superior to that of commercial sources and could become a limiting factor if used for qubit operation [23]. Further studies are needed to determine the phase noise requirements for qubit operation. Moreover, for qubit operation, the frequency of 140 MHz is too low. For operations involving electron spin states, GHz-range frequencies would be required. To achieve oscillation at higher frequencies, it is essential to reduce the parasitic capacitance of the TD and the circuit board. The parasitic capacitance of tunnel diodes exhibits significant variation, which is unlikely to be solely due to the pn-junction capacitance. Instead, this variation may originate from the diode packaging. Therefore, rather than relying on commercially available tunnel diodes, designing a tunnel diode specifically optimized for qubit readout or operation applications would be more effective. Minimizing the parasitic capacitance on the circuit board is also important. One possible approach is to fabricate the entire circuit board using microfabrication techniques. Another approach is to replace bulky varactor diodes with variable capacitors made from ferroelectric materials, such as strontium titanate (STO) [37]. Ferroelectric materials like STO not only allow for variable capacitance but also eliminate leakage currents, reducing circuit losses and providing a significant advantage over traditional varactor diodes.
@@
Considering that the cooling power of a typical cryogenic refrigerator’s 4 K stage is approximately 200 mW, and its MC stage provides about $400\,\mu\text{W}$ at 100 mK, it is feasible to place up to 20,000 microwave sources on the 4 K stage and 400 on the MC stage. These numbers are promising for realizing a scalable quantum computer; nevertheless, further improvements are necessary. In principle, lower power consumption of tunnel diodes can be fabricated by making the surface area of the tunnel diode smaller. While this would also lead to a reduction in output power, it can be compensated by adjusting the ratio of the number of turns in the spiral inductor to the number of turns in the pickup coil. However, the impact of back action on the oscillator must be carefully considered to ensure reliable operation. Furthermore, considering the need to apply a magnetic field for electron spin qubits, improvements such as using magnetic field-resilient superconducting materials like NbTiN instead of Nb and designing spiral inductors with these materials could be explored.
@@ 7 Conclusion
In conclusion, we developed a 140 MHz tunnel diode oscillator with low power consumption of 1 $\mu$W and characterized its performance on the 10 mK stage. The achieved output power and frequency fall within the range typically used for qubit readout of electrons in semiconductors used as qubits. The amplitude stability is better than that of the commercial microwave sources measured in this study, which is encouraging for its application in qubit readout.
@@ Acknowledgements
Appendix A Individual differences in BD-6 tunnel diodes
Fig. 6 compares the I-V curves of different BD-6 tunnel diodes measured at 25 K and RT. As stated in the BD-6 data sheet, the peak current at RT is 20 $\mu$A, which is consistent across all BD-6 diodes. However, at 25 K, significant variations among individual devices become apparent. Notably, BD-6 #02 does not exhibit negative resistance at low temperatures, meaning that it cannot be used for the cryogenic microwave source. The BD-6 diode used in this study exhibits a low-temperature I-V characteristic similar to that of BD-6 #03 but is a different individual device. Typically, a temperature of 25 K is low enough to evaluate and select a suitable BD-6 tunnel diode.
Figure 6: I-V curves of three different BD-6 tunnel diodes measured at RT and 25 K. The dots represent measurement points, and the lines connect these points to guide the eye. This measurement was performed using bare tunnel diodes, not those connected to an LC resonator for oscillation. As a result, no oscillation occurs, and thus no hysteresis is observed, unlike in Fig. 2(d)
Appendix B Temperature dependence
@@
This circuit is thermally anchored to the 10 mK stage, or the Mixing Chamber (MC) stage, and the MC stage temperature (MCT) was controlled while measuring a thermometer that is also thermally anchored to the MC stage. The sensitivity was found to be 0.3 kHz/mK in the MCT range of 60 mK to 120 mK, with no detectable dependence below 60 mK, aside from a slight hysteresis. This result is significantly better than the temperature dependence reported in previous studies above 1 K [8].
The observed temperature dependence between 60 mK and 120 mK is unlikely to be due to changes in the tunnel diode’s I-V curve, as no significant variation in its characteristics was observed in this range. Therefore, it is improbable that frequency shifts result from a change in the operating point. In prior studies measuring above 1 K [8], temperature dependence was attributed to the surface impedance, thermal expansion, and magnetic susceptibility induced by iron impurities in the copper used in the inductor. In contrast, this work uses a superconducting material, Nb, with a critical temperature $T_{c}$ of 6 K, in the inductor. Given that the measured temperature range is significantly below $T_{c}$, it is unlikely that the temperature dependence is primarily due to changes in the properties of Nb, though some contributions cannot be ruled out. Alternatively, the temperature dependence might arise from changes in the physical properties of the metals used in the PCB substrate.
Moreover, we note that we did not directly measure the temperature of the circuit itself, and its actual temperature may differ from MCT. The absence of detectable temperature dependence below 60 mK may be due to the fact that the diode’s temperature is likely saturated by the liberated heat and does not cool further. Please also note that the TDO circuit described in the main text is similarly thermally anchored to the MC stage, and its temperature was measured using a thermometer separately thermally anchored to the MC stage. However, while the qubit must be cooled to 10 mK, the TDO circuit does not necessarily need to be at 10 mK. For the purposes discussed in this work, this is not a concern.
Appendix C Varactor diode IV curve
@@
Appendix D DC voltage source stability
@@
Appendix E Time-domain data
@@
Figure 9: Time-domain data measured via AC port II using a 20 GS/s sampling rate oscilloscope (LeCroy WaveRunner 9054). The data include signals from the TDO powered by Yokogawa GS200 DC source (green) and a lead-acid battery (light green), as well as four different commercial microwave sources: Rigol DSA815 (black), Analog Devices ADF4351 (blue), and Vaunix Lab-Brick LMS-451D (red). The respective carrier frequencies are $f_{0}=$141.800 MHz, 141.841 MHz, 141.802 MHz, 141.806 MHz, and 141.803 MHz. For the TDO, $V_{\mathrm{TD}}=0.18$ V, and $V_{\mathrm{VD}}=-0.5$ V. For a consistency check, a synthetic sine signal at 141.8 MHz with 15 mV amplitude is also plotted. The raw signal is shown in (a), while (b) presents the signal after applying a digital bandpass filter with a 50 MHz bandwidth centered on the carrier frequency.
Appendix F Power spectra
@@
where the time-domain data $x_{n}$ is shown in Fig. 9(b) and $X_{k}$ represents the Fourier coefficients. The root-mean-square amplitude in volts was calculated as: